Opened 7 years ago

Closed 4 years ago

Last modified 15 months ago

#450 closed flight processing (fixed)

RG12/10, flight day 208/2012, Milton Keynes

Reported by: knpa Owned by: adbe
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/RG12_10-2012_208_Milton_Keynes

Data arrived from ARSF via SATA disk on 27/07/2012

RG12/10 Scientific objective:

RG12/10 Priority: 8h

PI: Karen Anderson

Sensors:

Camera (11/09/2012)
Eagle (13/11/2012)
Hawk (13/11/2012)
LiDAR (13/11/2012)
FW LiDAR (30/01/2013)

Change History (39)

comment:1 Changed 7 years ago by knpa

All the webcam files are blank for this project - will e-mail Phil about it.

Files for first eagle line were incorrectly named as datacube.*, have fixed.

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)

comment:3 Changed 7 years ago by adbe

Navigation Processed

comment:4 Changed 7 years ago by adbe

Camera delivery created

comment:5 Changed 7 years ago by adbe

Hyperspectral delivery created

comment:6 Changed 7 years ago by knpa

Delivery checking camera.

comment:7 Changed 7 years ago by knpa

Delivery check complete. No problems found. Ready to deliver.

comment:8 Changed 7 years ago by knpa

Hyperspectral delivery is actually just Eagle as we don't have a boresight for Hawk yet.

Beginning Eagle delivery check.

comment:9 Changed 7 years ago by knpa

  • Owner set to adbe
  • Status changed from new to assigned

comment:10 Changed 7 years ago by knpa

Finished delivery check.

All looks good apart from:

  • rename delivery to RG12_10-208-eagle-20120827 (do this for 206b too)
  • OSTNO2 file - think this should be in a misc directory. Don't forget to update ReadMe contents.
  • Remove Hawk examples from ReadMe Contents
  • In the filenumber->flightline name list, explain missing 046a or omit from this list. Get rid of blank Hawk section of this list.
  • In quality/issues section, explain better what is wrong with line 10, layman will not understand 'sct'

Have ran zip_mapped.sh

comment:11 Changed 7 years ago by adbe

Removed OSTN02_NTV2.gsb file
Removed Hawk example from ReadMe contents
Renamed to eagle delivery
Removed 046a, and Hawk column
Rewritten quality issues

comment:12 Changed 7 years ago by adbe

  • Description modified (diff)

Delivered Camera to Karen Anderson on 11/09/2012 on disk 44

comment:13 Changed 7 years ago by besm

Decided on pitch and roll errors for this flight, although I'm still checking them. To 4 significant figures, they are:

Pitch: -0.002870
Roll: +0.0001842

comment:14 Changed 7 years ago by besm

After checking the flightlines, most were pretty good, except the milton_x line (the one that crosses all other lines) for some reason. It was off by about 30cm on the ground. It seemed to be nearly completely pitch error, so the following values are what I've chosen for just the milton_x flight line:

Pitch: -0.003019
Roll: +0.0001842

comment:15 Changed 7 years ago by anhi

LiDAR delivery check:

  • Maximum and minimum Y values in "per flight line statistics" incorrect. Just remake tex file once lasinfo_reader.py is updated.
  • Looks like not all overlapping lines have an elevation difference recorded.
  • Remove the .tex, .out, .aux and .log Read_Me files from the delivery directory.

comment:16 Changed 7 years ago by besm

Done points 1 and 3. We discussed this and decided the case for the second point wasn't strong enough considering the accuracy of the resultant table anyway, so that's being left as-is.

Otherwise, should be ready to deliver. Marking as ready to delivery check.

comment:17 Changed 7 years ago by anhi

All looks fine now, ready to deliver without vectors.

comment:18 Changed 6 years ago by besm

Finally completed delivery for hawk, and ammended the eagle readme. Ready to delivery check. A couple of things to keep in mind are:

  • There are seperate deliveries for the hawk and eagle.
  • The script(s?) assumed there were 22 flightlines available when writing the readme, but there is no hyperspectral data for flightline 046b. The latex files have been edited by hand to correctly number the flightlines. I also wrote the correct mapping into a file called FLIGHTLINE_TO_FILE_MAPPING.txt in the processing directory and made it read-only.
  • Several different scripts fell down making this delivery, so the delivery checker should look carefully for things not being where they should be. I've already tried to make sure everything's there though, so shouldn't be any problems.

comment:19 Changed 6 years ago by anhi

Hawk delivery check complete:

  • Mention dropped scan in hawk line 14.
  • Hawk delivery readme lists eagle lines in the delivery contents which wont be there.

Proj_tidy.sh does not work with half deliveries so compared everything to the file structure and naming list manually

comment:20 Changed 6 years ago by besm

All fixed and ready to go.

comment:21 Changed 6 years ago by besm

Scratch that last comment. Delivery needs FW LiDAR. Awaiting FW boresight.

comment:22 Changed 6 years ago by besm

Sending without FW LiDAR after updating data quality report to the one found in ~arsf/documents_in_svn/lidar/reports/*1109.pdf

comment:23 Changed 6 years ago by besm

  • Description modified (diff)

comment:24 Changed 6 years ago by besm

Delivered everything, camera (again), hawk, eagle, LiDAR to Karen Anderson on 13/11/2012 on disk 133.

comment:25 Changed 6 years ago by besm

Completed full-waveform LiDAR delivery, ready for delivery checking.

comment:26 Changed 6 years ago by anhi

Full-Waveform delivery check:

No Problems

comment:27 Changed 6 years ago by besm

Delivered LiDAR full-waveform to Karen Anderson+Antoine Cottin on 30/01/2013 on either disk 148 or 150.

comment:28 Changed 6 years ago by besm

  • Description modified (diff)

comment:29 Changed 6 years ago by mark1

Deleted web-cam images as they are all corrupt 0 bytes

comment:30 Changed 5 years ago by emca

Preparing for archive

comment:31 Changed 5 years ago by emca

  • Component changed from Processing: general to Archiving

comment:32 Changed 5 years ago by emca

Eagle and Hawk delivered separately

Creating a joint delivery for archiving

comment:33 Changed 5 years ago by emca

Created the sct sync file (note: I got these values from the logfiles).

Sct values are as follows:

Flightline Eagle SCT Hawk SCT
1 -0.02 -0.05
2 -0.04 -0.05
3 -0.02 -0.05
4 -0.05 -0.04
5 -0.02 -0.05
6 -0.02 -0.06
7 -0.02 -0.06
8 -0.02 -0.06
9 -0.02 -0.06
10 -0.07 -0.06
11 -0.04 -0.06
12 -0.05 -0.06
13 -0.04 -0.06
14 -0.02 -0.06
15 -0.02 -0.06
16 -0.02 -0.06
17 -0.02 -0.07
18 -0.04 -0.04
19 -0.02 -0.04
20 -0.07 -0.06
21 -0.02 -0.02

comment:34 Changed 5 years ago by emca

Added the SCT values to the hyperspectral processing config file

comment:35 Changed 5 years ago by emca

Painstakingly created a joint hyperspectral delivery.

Project now ready for archive.

comment:36 Changed 5 years ago by emca

Email sent to Wendy 04/12/2013

comment:37 Changed 4 years ago by dac

Reprocessed LAS files in ALSPP with AGC values. Sent to Steve Hancock via FTP 11/02/2014

comment:38 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Archived data already available at NEODC (without AGC values in FW data). Have uploaded most recent delivery.

comment:39 Changed 15 months ago by asm

Archiving

Reprocessed LiDAR data has been uploaded. JSON files created and uploaded as well and CEDA has been notified (31/01/2018).

Note: See TracTickets for help on using tickets.